# AnesSuite: A Comprehensive Benchmark and Dataset Suite for Anesthesiology Reasoning in LLMs

### v3

@@ G.1 Evaluation on Out of Distribution Benchmarks
To further assess the generalization capability of the model trained on anesthesiology reasoning data, we evaluated Morpheus across three general medical benchmarks (MedQA, MedMCQA, and MedExpQA; Jin et al., 2021; Pal et al., 2022; Alonso et al., 2024) and three general-domain benchmarks (MMLU, MMLU-Pro, and AGIeval; Wang et al., 2023a; 2024b; Zhong et al., 2024). Consistent with our previous experiments, we employed the same framework used in AnesBench to extract answers for multiple-choice questions. Notably, for AGIeval, we restricted the evaluation to its 18 MCQ tasks, whereas comprehensive evaluations were conducted for all other benchmarks. The medical subset of MMLU includes the tasks anatomy, clinical_knowledge, college_biology, college_medicine, high_school_biology, medical_genetics, nutrition, professional_medicine, and virology.
@@ G.2 Self-Consistency in Test Time Scaling
G.3 Evaluation on the Open-ended Subset
@@
The results presented in Table 23 reveal a strong correlation between the LLM-based scores (GPT and Gemini) and the original MCQ accuracy on AnesBench, which validates the use of MCQs as a reliable proxy for complex reasoning. In contrast, automated metrics like BLEU and F1 scores were consistently low (mostly below 0.25), underscoring their inadequacy for assessing the nuanced, multi-step logic required in medical decision-making. Therefore, we prioritize MCQs in our benchmark as a more robust and standardized method for evaluating reasoning performance.
G.4 Evaluation on CAB
@@ H.3 Collection Process
A4: The authors of this paper.
@@
H.4 Preprocessing/cleaning/labeling
@@ 1. Who will be supporting/hosting/maintaining the dataset?
A2: The public contact email will be posted post-acceptance.

### v4

@@ G.1 Evaluation on Out of Distribution Benchmarks
To further assess the generalization capability of the model trained on anesthesiology reasoning data, we evaluated Morpheus across three general medical benchmarks (MedQA, MedMCQA, and MedExpQA; Jin et al., 2021; Pal et al., 2022; Alonso et al., 2024) and three general domain benchmarks (MMLU, MMLU-Pro, and AGIeval; Wang et al., 2023a; 2024b; Zhong et al., 2024). Consistent with our previous experiments, we employed the same framework used in AnesBench to extract answers for multiple-choice questions. Notably, for AGIeval, we restricted the evaluation to its 18 MCQ tasks, whereas comprehensive evaluations were conducted for all other benchmarks. The medical subset of MMLU includes the tasks anatomy, clinical_knowledge, college_biology, college_medicine, high_school_biology, medical_genetics, nutrition, professional_medicine, and virology.
@@
Table 24: Analysis results of the model’s reasoning hallucinations on the AnesBench-English system2 subset. The numbers in the cells respectively represent the quantity of various types of reasoning sentences in the sampled subset and their proportion of the total reasoning sentences.
Model Valid NS CR OE MC
AnesBench-English System2
Morpheus-7B 953 (76.2%) 206 (16.5%) 13 (1.0%) 66 (5.3%) 12 (1.0%)
Qwen2.5-7B-Instruct 181 (68.8%) 55 (20.9%) 1 (0.4%) 18 (6.8%) 8 (3.0%)
Morpheus-14B 1119 (81.4%) 191 (13.9%) 5 (0.4%) 51 (3.7%) 8 (0.6%)
Qwen2.5-14B-Instruct 254 (77.0%) 51 (15.5%) 4 (1.2%) 18 (5.5%) 3 (0.9%)
Morpheus-32B 1179 (83.6%) 151 (10.7%) 4 (0.3%) 62 (4.4%) 14 (1.0%)
Qwen2.5-32B-Instruct 233 (78.2%) 43 (14.4%) 1 (0.3%) 15 (5.0%) 6 (2.0%)
Qwen2.5-72B-Instruct 230 (74.9%) 52 (16.9%) 5 (1.6%) 13 (4.2%) 7 (2.3%)
Notes: NS: Non-Sequitur, CR: Causal Reversal, OE: Over-extrapolation, MC: Medical Contraindication.
@@ G.2 Self-Consistency in Test Time Scaling
G.3 Hallucination Evaluation
Beyond the accuracy on MCQs, ensuring the accuracy of each reasoning step generated by the model is of greater importance when translating research findings into practical clinical applications. Therefore, we performed a detailed hallucination evaluation on the outputs of the several models presented in Tab. 4, utilizing an LLM-as-Judge paradigm.
Specifically, our analytical pipeline comprises several steps. First, we sampled 50 questions from System 2 of the English subsets of AnesBench. For these questions, we employed spaCy 444spaCy: https://spacy.io/ to segment each model’s response into independent sentences. Subsequently, the context and each individual sentence were fed into GPT-OSS-120B (Agarwal et al., 2025) for coreference resolution, while simultaneously filtering for complete sentences that contained explicit reasoning behaviors. Finally, GPT-OSS-120B, together with an evaluation prompt, was used to assess each sentence for the presence of four common reasoning-hallucination patterns during inference: Non-Sequitur, Causal Reversal, Medical Contraindication, and Over-extrapolation.
As shown in Tab. 24, the results indicate that among these hallucination types, Non-Sequitur and Over-extrapolation are the most prevalent. This suggests that models are highly susceptible to broken logical chains or drawing conclusions based on insufficient evidence, highlighting the need for future research to employ methods such as process reward model (Lightman et al., 2023) and retrieval-augmented generation (Lewis et al., 2020; Jiang et al., 2025) to enhance reasoning coherence and mitigate issues of insufficient evidence. Moreover, scaling these retrieval systems to accommodate large and multimodal clinical data will be increasingly critical (Müller & Unay, 2017; Shen et al., 2022; 2024b; 2024a).
Comparing the models, a larger parameter scale generally correlates with stronger reasoning validity in most cases. A direct comparison between the Morpheus and Qwen2.5 models reveals that Morpheus tends to generate more extensive reasoning steps. Notably, Morpheus demonstrated a consistent improvement in reasoning validity over the baseline Qwen2.5, thereby substantiating the effectiveness of our training data and methodology. Furthermore, across various parameter scales, the Morpheus models consistently maintain the MC (high-risk hallucination) error rate at or below 1.0%, which likewise demonstrates the enhanced safety profile of the models trained with our resources.
Prompt used for Hallucination Evaluation SYSTEM PROMPT You are an expert Clinical Logic Judge. Your task is to evaluate the logical soundness of a single reasoning step generated by a medical AI. You are NOT verifying isolated facts; you are judging the causal link and logical deduction within the sentence. Does the stated premise or fact logically justify the clinical conclusion or recommendation? # EVALUATION CRITERIA: Evaluate the internal logic of the Target_Sentence for the following fatal flaws: - Non-Sequitur: The conclusion does not logically follow from the premise (e.g., "Patient has a headache, therefore amputate the leg"). - Causal Reversal: Mixing up cause and effect. - Medical Contraindication: Recommending a treatment that is logically contradicted by the premise stated in the very same sentence (e.g., "Given the severe bleeding, administer heparin"). - Over-extrapolation: Jumping to a definitive, extreme conclusion without sufficient logical bridging. # EVALUATION CATEGORIES: 1. "Valid": The reasoning is medically sound, coherent, and the conclusion logically follows the premises provided in the sentence. 2. "Flawed": The reasoning contains a logical fallacy, contradiction, or medically dangerous leap in logic. # INSTRUCTIONS: Analyze the "Target_Sentence". Output ONLY a valid JSON object with the following structure: { "logic_judgment": "<Must be exactly one of: Valid, Flawed>", "flaw_type": "<If Flawed, briefly name the issue (e.g., Non-Sequitur, Contradiction). If Valid, output null>", "rationale": "Briefly explain why the logical bridge holds up or breaks down." } # EXAMPLES: Target_Sentence: "Because the patient is allergic to penicillin, we should prescribe amoxicillin." Output: {"logic_judgment": "Flawed", "flaw_type": "Contradiction", "rationale": "Amoxicillin is a penicillin-class antibiotic, so the premise of allergy directly contradicts the treatment recommendation."} Target_Sentence: "Given the elevated troponin levels and chest pain, an ECG should be performed to evaluate for myocardial infarction." Output: {"logic_judgment": "Valid", "flaw_type": null, "rationale": "Elevated troponin and chest pain are classic signs of cardiac injury, making an ECG a logical next diagnostic step."}
G.4 Evaluation on the Open-ended Subset
@@
The results presented in Tab. 25 reveal a strong correlation between the LLM-based scores (GPT and Gemini) and the original MCQ accuracy on AnesBench, which validates the use of MCQs as a reliable proxy for complex reasoning. In contrast, automated metrics like BLEU and F1 scores were consistently low (mostly below 0.25), underscoring their inadequacy for assessing the nuanced, multi-step logic required in medical decision-making. Therefore, we prioritize MCQs in our benchmark as a more robust and standardized method for evaluating reasoning performance.
@@
G.5 Evaluation on CAB
@@ 2. How many instances are there in total (of each type, if appropriate)?
A4: The data collection was primarily conducted by the lead authors of this paper. Additionally, Tao Huang and Zhiwei Wang from the School of Computer Science, Wuhan University, assisted in the data collection process.
@@
Preprocessing/cleaning/labeling
@@ 1. Who will be supporting/hosting/maintaining the dataset?
A2: Users can contact the authors via the email address on the first page of the main text.